OpenStudy (anonymous):

if the sum of three consecutive integers is 987. find the integers

OpenStudy (amistre64):

987/3 = 325 something closer to that one id assume

OpenStudy (anonymous):

x+x+1+x+2=987 3x+3=987 3x=987-3 x=984/3=328 Three consecutive integers are 328,329,330

OpenStudy (anonymous):

thank you

OpenStudy (pokemon23):

HERO !